OpenStudy (anonymous):

if we have f(x)=1+x divided by 1-x, can the f(x+h)=1+x+h divided by 1-x-h

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

1-(x) when x = x+h ... we have 1-(x+h) = 1-x-h

OpenStudy (amistre64):

i personally feel it a bit confusing at first to be replacing x parts with x parts .... i feel this would be much more intuitive as: f(x) = (f(a) - f(a+h))/h

OpenStudy (amistre64):

lol, got my fs mixed around

OpenStudy (anonymous):

yep it has to be solved using that definition